OpenStudy (anonymous):

can someone please explain to me how to simplify the expression below (5x)^2

OpenStudy (kinggeorge):

You can essentially distribute the exponent to the values in the parentheses. \[(5x)^2=(5^2)(x^2)=25x^2\]

OpenStudy (anonymous):

i don't get it?

OpenStudy (kinggeorge):

In your original equation, you're squaring the product of two things. Instead, you can square each thing first, and then multiply them together.

OpenStudy (anonymous):

Just like (a+b)*c = ac + bc, exponents work the same way. (ab)^c = a^c*b^c

OpenStudy (kinggeorge):

Perhaps a more clear way to explain it would be \[(5x)^2=(5x)(5x)=(5\cdot 5)(x \cdot x)=25x^2\]